”The length of a rectangular garden is three times the width. What are
the dimensions of the garden, if the perimeter is 32m?”

Answer by ikleyn(52788)

.
Let W be the width.


Then the length is 3W.


Perimeter = W + 3W + W + 3W = 32


8W = 32  =======>  W = 32/8 = 4.


ANSWER.  The dimensions are  4 meters and 3*4 = 12 meters.

Solved.